OpenStudy (anonymous):

find two solutions to the equation 3x+5y=10. write the solutions as ordered pairs.

OpenStudy (anonymous):

Please help!!

OpenStudy (akashdeepdeb):

3x + 5y = 10 As this is a line which MAY look like this, there are INFINITE number of (x,y) combinations, satisfying the line. 

OpenStudy (akashdeepdeb):

To find any 2 ordered pairs. Take any simple x or y value and find the corresponding y or x value. For eg. If x = 1 then 3x + 5y = 10 3(1) + 5y = 10 = 5y = 7 y = 7/5 You may find many such points, just like that! :D

OpenStudy (texaschic101):

exactly...just sub a number in for x....any number, and then find y. That is an ordered pair that is a solution to your equation.